# Score-based diffusion nowcasting of GOES imagery

### v1

@@ Abstract
Clouds and precipitation are important for understanding weather and climate. Simulating clouds and precipitation with traditional numerical weather prediction is challenging because of the sub-grid parameterizations required. Machine learning has been explored for forecasting clouds and precipitation, but early machine learning methods often created blurry forecasts. In this paper we explore a newer method, named score-based diffusion, to nowcast (zero to three hour forecast) clouds and precipitation. We discuss the background and intuition of score-based diffusion models - thus providing a starting point for the community - while exploring the methodology’s use for nowcasting geostationary infrared imagery. We experiment with three main types of diffusion models: a standard score-based diffusion model (Diff); a residual correction diffusion model (CorrDiff); and a latent diffusion model (LDM). Our results show that the diffusion models are able to not only advect existing clouds, but also generate and decay clouds, including convective initiation. These results are surprising because the forecasts are initiated with only the past 20 mins of infrared satellite imagery. A case study qualitatively shows the preservation of high resolution features longer into the forecast than a conventional mean-squared error trained U-Net. The best of the three diffusion models tested was the CorrDiff approach, outperforming all other diffusion models, the traditional U-Net, and a persistence forecast by one to two kelvin on root mean squared error. The diffusion models also enable out-of-the-box ensemble generation, which shows skillful calibration, with the spread of the ensemble correlating well to the error.
@@
Knowing where clouds and precipitation will be is important for everything from weather forecasts to telecommunications. Unfortunately, simulating clouds and precipitation is a challenge for physics-based approaches. Machine learning methodology can help, but older methodologies often result in blurry forecasts. In this paper, we explore a newer machine learning method from the computer science literature, named diffusion, which enables clearer and more realistic looking forecasts. We specifically look at how the diffusion models work and provide some base intuition to accelerate others using diffusion in their Earth and Environmental Science research. We explore how these diffusion models work for forecasting satellite infrared imagery. Our results show skillful performance of these new diffusion models, outperforming prior machine learning techniques.
@@ 1 Introduction
Clouds and precipitation play a vital role in understanding the radiative budget of Earth and understanding the past, current, and future climate. Furthermore, the exact location and timing of clouds and precipitation have natural applications to forecast weather, solar power generation (Zhang et al., 2018) and laser communications (Craddock et al., 2024). The main methods for forecasting clouds and precipitation are Lagrangian advection schemes (e.g., Bowler et al., 2006; Pulkkinen et al., 2019), traditional numerical weather prediction (e.g., Griffin et al., 2017, 2024), optical flow (e.g., Shakya and Kumar, 2019), or machine learning (e.g., Berthomier et al., 2020; Kellerhals et al., 2022).
The use of machine learning in meteorology is not new and has been growing rapidly (c.f., Figure 1 in Chase et al., 2022). In fact, neural networks have shown skill in forecasting clouds and precipitation (e.g., Shi et al., 2015; Caseri et al., 2022; Andrychowicz et al., 2023). The challenge with the early types of neural networks, such as convolutional neural networks, is that while they are skillful they often result in blurry predictions (Bouallègue et al., 2024) which may be less useful to forecasters. One way to combat the blurry issue has been the use of Generative AI methods, such as Generative Adversarial Networks (GANs; Goodfellow et al., 2014; Andrianakos et al., 2019; Ravuri et al., 2021). Unfortunately, GANs are notorious for being difficult to train (Saxena and Cao, 2021) and have suboptimal uncertainty quantification (e.g., underdispersive ensemble forecasts; Harris et al., 2022; Price and Rasp, 2022). A newer Generative AI method to combat blurry output is known as diffusion models (Ho et al., 2020; Song et al., 2021; Karras et al., 2022) which has already been used in cloud and precipitation forecasting (Hatanaka et al., 2023; Zhang et al., 2023; Dai et al., 2024; Nai et al., 2024; Wang et al., 2024; Yu et al., 2024) and, more generally, in data-driven weather forecasting (Manshausen et al., 2024; Pathak et al., 2024; Price et al., 2025). Diffusion models use machine learning to iteratively solve a carefully constructed differential equation (inspired by those governing the physical process of heat diffusion, thus the name) to turn random noise into generated images that are consistent with the distribution of the data they are trained on. More casually speaking, that look like they could have come from the training set.
The goal of this paper is to specifically explore score-based diffusion models (Song et al., 2021; Karras et al., 2022) and their ability to nowcast infrared brightness temperatures of geostationary satellites. The end goal of such a nowcasting model is eventually to forecast clouds and precipitation. In the exploration of this relatively new methodology, we discuss the hurdles encountered in implementing score-based diffusion models with Earth and Environmental Science data. Furthermore, we also discuss some often overlooked steps needed for meteorological applications that are not addressed in the computer science literature on diffusion models. The goal of this added discussion and open-access code is to accelerate the use of score-based diffusion models among domain scientists in the Earth and Environmental Sciences.
In our exploration of score-based diffusion models we try three types that all follow the main design approach from Karras et al. (2022), also known as Elucidating the Design space of diffusion-based generative Models (EDM). The three specific types are
@@
    A plain diffusion model (Diff; Karras et al., 2022),
@@
    A residual correction diffusion model (CorrDiff; Mardani et al., 2025; Pathak et al., 2024),
@@
    A latent diffusion model (LDM; Rombach et al., 2022; Leinonen et al., 2023).
This paper is organized as follows. Section 2 contains the overview of score-based diffusion, including a brief discussion of the mathematical foundations of diffusion models. Section 3 provides background on the geostationary dataset we use, the steps to prepare it for using score-based diffusion, how it is adapted for the satellite forecasting task, what specific types of machine learning models are trained, and some exploration of pre-trained autoencoders. Section 4 shows the results of several score-based diffusion models trained to forecast satellite imagery and discusses their performance compared to a couple of baseline forecasts, persistence and a mean squared error trained U-Net. Lastly, Section 5 summarizes the paper and provides an outlook of improvements that could be made.
2 Background on Score-Based Diffusion Models
@@ 2.1 Preface
All diffusion models discussed in this paper follow a score-based diffusion framework (EDM; Song et al., 2021; Karras et al., 2022). Before we dive into the details of the EDM framework we acknowledge another common framework for diffusion models, named Denoising Diffusion Probabilistic Models (DDPM; Ho et al., 2020). We initially explored DDPMs for this project given their wide use in the literature (e.g., Asperti et al., 2024; Chen et al., 2023; Gao et al., 2023; Srivastava et al., 2024; Zhong et al., 2024), but realized the computational demand would not be viable given our hardware and the temporal demand of an operational nowcasting product. Instead, after seeing the results from Li et al. (2024), Mardani et al. (2025), Pathak et al. (2024), and Price et al. (2025), we decided to pursue score-based diffusion models which we found easier to train and evaluate. The EDM framework can be interpreted as a generalization of the DDPM approach. For more information on this topic, see Ho et al. (2020) for DDPMs, see Song et al. (2021) and Karras et al. (2022) for the connection between the DDPM and EDM framework, and see references therein and several Earth and Environmental Science papers using DDPMs (Asperti et al., 2024; Chen et al., 2023; Gao et al., 2023; Srivastava et al., 2024; Zhong et al., 2024).
@@ 2.2 Basic idea of Diffusion models
Figure 1: Conceptual overview of diffusion model approach for unconditional image generation with stochastic differential equations, using Equations (1) and (2) from Song et al. (2021) for the (a) forward and (b) backward process. See Section 22.3 for more explanation. Note that this is an idealized representation, as the reconstructed image generally does not exactly match the original image, because the backward process is imperfect.
Generally, diffusion models can be thought of as tools to learn, and later sample from, data distributions. A diffusion model is trained on a large set of representative samples of the distribution, typically images. For example, if a diffusion model is trained on an extensive set of sample images from a specific satellite channel, the diffusion model learns to approximate this distribution. Once trained the diffusion model can provide new samples from the distribution, i.e., it generates new (approximate) satellite imagery for that channel. This type of random but realistic image generation is called unconditional image generation. Although such ability is useful for many applications, we are more interested here in the ability to generate specific imagery, known as conditional image generation. Conditional image generation would allow us, for example, to use information from a different channel of imagery or to forecast satellite imagery 10 minutes from now by conditioning on relevant imagery. In Sections 22.2 to 22.5 we focus entirely on diffusion models for unconditional imagery, while Section 22.6 discusses how to add the conditioning.
To perform diffusion image generation, these models utilize a mathematically well-described diffusion process for both the training of the underlying neural network and generating new images. Namely, to learn the distribution from a given data set of images (training set), training images are partially destroyed by applying a diffusion process to the image (e.g., adding noise; Fig. 1a). The diffusion model is then trained to restore the structure of the images from their degenerated versions (Fig. 1b). This process teaches the model to generate imagery of the same structure as the images in the training data set. In other words, the model learns to approximate the distribution represented by the training data set. Diffusion models are thus based on developing a model to reverse the diffusion equation, an approach developed mathematically over 40 years ago by Anderson (1982). However, certain components of this approach are hard to implement in practice, and thus the full potential of this approach was made possible only once those components could be implemented through neural networks, as proposed by Sohl-Dickstein et al. (2015).
@@ 2.3 Score-based Diffusion - Mathematical Foundation
The equations discussed here are all obtained from Song et al. (2021) and Karras et al. (2022); however, we attempt to provide additional intuition behind the equations in those papers. The key aspect of diffusion models is that they leverage well-described stochastic differential equations to transform images through a diffusion process (Fig. 1), which in this application is the addition of random noise. Physical diffusion describes the effects of the random motion of particles towards a state of maximal entropy; in our case we are using the same mathematical principles to “diffuse” the original image by applying incremental random changes to the pixel values such that eventually, those random changes completely overwhelm the original signal that was present.
During the forward process, noise is added to the original imagery, while during the backward process noise is removed to restore image structure. For ML applications using diffusion models, noise is added and then removed from original imagery during training and the model learns weights associated with this process, so that during inference the ML model can then generate an image from pure random noise. To describe this process, consider the stochastic differential equation for the forward process of diffusion (i.e., the destruction of an image by the addition of noise):
$\mathrm{d}\bm{x}=\bm{f}(\bm{x},t)\;\mathrm{d}t+{g}(t)\;\mathrm{d}\bm{w}.$ (1)
In Equation (1), $\bm{x}$ is an image that is being transformed by adding random noise, $\bm{f}(\bm{x},t)$ is known as the drift coefficient, ${g}(t)$ is the diffusion coefficient, and $\bm{w}$ is the standard Wiener process (e.g., random motion, Brownian motion). The $t$ here is referred to as time and to be clear this is diffusion time where increased “time” represents increased noise in the image rather than a temporal relationship. Because we are using a differential equation, the reverse process can be derived (i.e., for our application a structured image can be generated from random noise). This is known as the backwards process and is derived in Song et al. (2021):
$\mathrm{d}\bm{x}=[\;\bm{f}(\bm{x},t)-g(t)^{2}\;\nabla_{\bm{x}}\log p_{t}(\bm{x})\;]\;\mathrm{d}t+g(t)\;\mathrm{d}\bar{\bm{w}},$ (2)
where $p_{t}(x)$ is the probability of a particular noisy image $\bm{x}$ under the distribution derived from the training data at noise level $t$, $\nabla_{\bm{x}}\log p_{t}(\bm{x})$ is known as the score function, and $\bar{\bm{w}}$ represents the standard Wiener process but in reverse.
As written, Equation (2) is abstract; however, Karras et al. (2022) reformulated the equation into more intuitive components using specific choices for $\bm{f}(\bm{x},t)$ and $g(t)$ that relate to parameters of a Gaussian noise distribution:
$\mathrm{d}\bm{x}_{\pm}=\underbrace{-\dot{\sigma}(t)\sigma(t)\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)\big{)}\,\mathrm{d}t}_{\text{probability flow ODE}}\,\\ \pm\,\underbrace{\underbrace{\beta(t)\sigma(t)^{2}\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)\big{)}\,\mathrm{d}t}_{\text{deterministic noise decay}}+\underbrace{\sqrt{2\beta(t)}\sigma(t)\,\mathrm{d}\omega_{t}}_{\text{noise injection}}}_{\text{Langevin diffusion SDE}}.$ (3)
Here $\sigma(t)$ is the standard deviation of the Gaussian distribution of noise added to the image at time $t$, $p(x,\sigma(t)$) is a reformulation of $p_{t}(x)$ in terms of $\sigma$, and $\beta(t)$ is a parameter describing the relative scale of additional noise re-added during denoising to make this a stochastic differential equation. The first term on the right of Equation (3) is the probability flow ordinary differential equation (ODE) portion, which is the primary vehicle to move an image towards or away from the desired data distribution. The remaining terms on the right of Equation (3) represent the Langevin diffusion stochastic differential equation (SDE) which modulates the probability flow ODE by adding and removing additional random noise from the image during both the forward and reverse processes.
For now, let us focus on just the probability flow ODE,
$\mathrm{d}\bm{x}=-\dot{\sigma}(t)\sigma(t)\underbrace{\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)}_{\text{score function}}\big{)}\,\mathrm{d}t.$ (4)
Section 22.5 will discuss adding back in the (SDE) portion of Equation (3). Casting Equation (4) into the satellite imagery forecasting task of this paper, $\bm{x}$ is the vector of noisy GOES IR images and $\mathrm{d}\bm{x}$ is then how to change the images $\bm{x}$ to add or remove noise. To remove noise, it is crucial to know the amount of noise in $\bm{x}$, which is described by the parameter $\sigma(t)$, where $\sigma$ represents the standard deviation of the Gaussian noise that was added to the image (i.e., standard deviation of the normal curve). The parameter $t$ is referred to as time in many papers, but to not get confused with the forecasting objective of this paper, view $t$ as a distance metric, where larger $t$ means we are farther from the original, non-noisy images. $\dot{\sigma}(t)$ is the first derivative of $\sigma(t)$ with respect to $t$. Karras et al. (2022) assumes $\sigma(t)$ is $t$, which makes $\dot{\sigma}(t)$ equal to one and simplifies Eq. (4) to $\mathrm{d}\bm{x}=-t\;{\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)}\big{)}\,\mathrm{d}t$. The last term to discuss is the score function, which is the key aspect of Equation (4). Note that the score function is the only unknown in Equation (4) and can be learned by a neural network.
Figure 2: Describing the diffusion process through the pixel-noise ($\mathbf{x}-\sigma$) diagram. (a) Using a single image (Hurricane Lee), the normalized histogram of the pixel values are shown in the vertical lines. Each line color represents different amounts of noise added to the image, where increasing noise goes from blue to green to orange and correspond to the same outline colored images. (b) The density of points (i.e., probability) of all images in the training database with stochastic amounts of noise added to them.
To help describe what the score function represents, consider our GOES IR data as a visual example (Fig. 2). Figure 2 shows the $\bm{x}-\sigma(t)$ diagram included in many score-based diffusion papers. This diagram shows a simplified, projected version of the high-dimensional space containing our data distribution at every noise level. This is useful to visualize particularly because solving the diffusion ODE or SDE gives a trajectory through this space that describes how a particular image was generated. Unlike many of the purely theoretical versions of this figure in the literature, the version shown in Fig. 2 is generated using the actual distribution of GOES imagery in the training set (Fig. 2b), and using example GOES images of Hurricane Lee with various levels of noise added to them (Fig. 2a). In this $\bm{x}-\sigma(t)$ space, we calculate a single histogram across all pixels of an image, i.e., we treat the values of all pixels as if they belonged to a single pixel, effectively collapsing the image dimension to one. Each noisy image can then be projected down to a single probability density function (PDF; i.e., normalized histogram of image values) for the corresponding exact value of $\sigma(t)$. The first image has no noise added to it, and the corresponding PDF is the black line. The other images have increasing values of $\sigma(t)$ with corresponding PDFs shown by the blue, green and orange lines. If our training dataset consisted of this single Hurricane Lee image and was comprised of completely independent pixels, then $p\big{(}\bm{x};\sigma(t)\big{)}$ in Equation (4) would be exactly these PDFs, just evaluated at more $\sigma(t)$ values. Extending this to our entire dataset of GOES IR images results in the color shading in Figure 2b, which is a 1-dimensional view of $\log p\big{(}\bm{x};\sigma(t)\big{)}$.
In practice, spatial context is very important for the reverse flow process and as such, these distributions are high-dimensional. We provide these one-dimensional views to illustrate the intuition of how diffusion works. The partial derivative of this probability space with respect to $\bm{x}$, namely $\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)\big{)}$, points towards regions of higher probability in the distribution at each noise level, i.e., it tells us how to change $\bm{x}$ to reach a higher probability image. It is not surprising, then, that to run the probability flow ODE backwards in $\sigma(t)$, this score function can be used to move an image towards higher-probability regions. You can view the score $\nabla_{\bm{x}}\log p\big{(}\bm{x};\sigma(t)\big{)}$ as the instantaneous direction we need to follow from a pure noise image (which would be located on the right y-axis in Figure 2) in order to generate a realistic image from the GOES IR dataset (left y-axis). Hopefully it is clear that if the score function could be explicitly known, then one could generate images from the training dataset from any random noise start. How do we get the score function? This is where neural networks come in.

### v2

# How to use score-based diffusion in earth system science: A satellite nowcasting example

@@ Abstract
Machine learning (ML) is used for many earth science applications; however, traditional ML methods trained with squared errors often create blurry forecasts. Diffusion models are an emerging generative ML technique with the ability to produce sharper, more realistic images by learning the underlying data distribution. Diffusion models are becoming more prevalent, yet adapting them for earth science applications can be challenging because most articles focus on theoretical aspects of the approach, rather than making the method widely accessible. This work illustrates score-based diffusion models with a well-known problem in atmospheric science: cloud nowcasting (zero-to-three-hour forecast). After discussing the background and intuition of score-based diffusion models using examples from geostationary satellite infrared imagery, we experiment with three types of diffusion models: a standard score-based diffusion model (Diff); a residual correction diffusion model (CorrDiff); and a latent diffusion model (LDM). Our results show that the diffusion models not only advect existing clouds, but also generate and decay clouds, including convective initiation. A case study qualitatively shows the preservation of high-resolution features longer into the forecast than a conventional U-Net. The best of the three diffusion models tested was the CorrDiff approach, outperforming all other diffusion models, the conventional U-Net, and persistence. The diffusion models also enable out-of-the-box ensemble generation with skillful calibration. By explaining and exploring diffusion models for a common problem and ending with lessons learned from adapting diffusion models for our task, this work provides a starting point for the community to utilize diffusion models for a variety of earth science applications.
@@
Machine learning is an invaluable tool for earth science applications, but they often result in blurry images, predictions, or forecasts. Diffusion models are an emerging technique to enable more realistic looking images. This work intuitively explains the diffusion modeling process and explores diffusion models for forecasting satellite imagery. Our results show skillful performance by diffusion models, outperforming traditional machine learning techniques. We discuss lessons learned from applying different diffusion models and trade-offs between performance and computing requirements that are important considerations for deploying these models. The main goal is to provide basic intuition for diffusion modeling and to pass along our hard-earned tips to help accelerate others towards using diffusion models for their own Earth and Environmental Science research tasks.
@@ 1 Introduction
The use of machine learning in meteorology has been growing rapidly (c.f., Figure 1 in Chase2022); however, the challenge with neural networks trained with mean squared error losses is that while they are skillful, they often result in blurry predictions (bouallegue2024rise). One way to combat the blurry issue has been the use of generative AI methods, such as Generative Adversarial Networks (GANs; Goodfellow2014; Andrianakos2019; Ravuri2021). Unfortunately, GANs are notorious for being difficult to train (Saxena2021) and have suboptimal uncertainty quantification (e.g., under-dispersive ensemble forecasts; Harris2022; Price2022). An emerging generative AI method to combat blurry output is diffusion modeling (ho2020; Song2021; Karras2022). Diffusion models use machine learning to iteratively solve a carefully constructed differential equation (inspired by the mathematical equations governing the physical process of heat diffusion, thus the name) to turn random noise into generated text or images that are consistent with the distribution of the data they are trained on.
Diffusion models have much potential for Earth System science applications; however, one challenge to their widespread use is that they build on complex mathematical concepts and there are few resources available for earth system scientists to easily learn those concepts. While numerous papers in the Earth System science literature use diffusion models, they generally refer readers to computer science literature to learn about diffusion models. Currently, the main resources to understand diffusion models are the original computer science papers (e.g., Song2021; Karras2022), which take considerable time to digest and to put into practice. The goal of this paper is thus to provide an accessible introduction to diffusion models which is tailored to domain scientists in atmospheric science in the hope of making diffusion models more accessible to this community. Our approach consists of:
@@
    Providing a self-contained, easy-to-understand introduction of the key foundational concepts of diffusion models. Specifically, we focus here on score-based diffusion models (Song2021; Karras2022), which to date are the most commonly used type of diffusion models in atmospheric science.
@@
    Illustrating the use of diffusion models for a well-known dataset and problem in atmospheric science. Namely, we explore their use to nowcast infrared brightness temperatures of geostationary satellites. The end goal of such a nowcasting model is eventually to forecast clouds and precipitation, but here it is used primarily for illustrative purposes.
@@
    Using the application of nowcasting satellite imagery to illustrate the numerous trade-offs and corresponding design decisions one has to navigate when using diffusion models.
In the exploration of this relatively new methodology, we discuss the hurdles encountered in implementing score-based diffusion models within the Earth System science community, including some often overlooked steps needed for meteorological applications that are not addressed in the computer science literature on diffusion models. The goal of this added discussion and open-access code is to accelerate the use of score-based diffusion models among domain scientists in the Earth System sciences.
The application we are utilizing to explore diffusion modeling is cloud and precipitation nowcasting. Clouds and precipitation play a vital role in understanding the radiative budget of Earth and understanding the past, current, and future climate. Furthermore, the exact location and timing of clouds and precipitation have natural applications to forecast weather, solar power generation (Zhang2018) and laser communications (Craddock2024). The main methods for forecasting clouds and precipitation are Lagrangian advection schemes (e.g., Bowler2006; Pulkkinen2019), traditional numerical weather prediction (e.g., Griffin2017; Griffin2024), and optical flow (e.g., Shakya2019). Recently, machine learning has been rapidly employed for forecasting clouds and precipitation using traditional deep learning techniques (e.g., Shi2015; Berthomier2020; Caseri2022; Kellerhals2022), generative adverserial models (e.g., Andrychowicz2023), and diffusion modeling (Hatanaka2023; Zhang2023; Dai2024; Nai2024; Wang2024; Yu2024). In addition to becoming prevalent in nowcasting, diffusion models are also being used more generally in data-driven weather forecasting (Couairon2024; Manshausen2024; Pathak2024; Price2025).
Here, we are tasking the model with forecasting geostationary satellite infrared (IR) imagery, focusing on nowcasting (ten minutes out to three hours) 10.3 $\mu m$ brightness temperatures. In our exploration of score-based diffusion models we try three types that all follow the main design approach from Karras2022, also known as Elucidating the Design space of diffusion-based generative Models (EDM). The three specific types are
  1. 1.
    A plain diffusion model (Diff; Karras2022),
  2. 2.
    A residual correction diffusion model (CorrDiff; Mardani2025; Pathak2024),
  3. 3.
    A latent diffusion model (LDM; Rombach2022; Leinonen2023).
This paper is organized as follows. Section 2 contains an intuitive overview of score-based diffusion, including a brief discussion of the mathematical foundations of diffusion models. Section 3 provides background on the geostationary satellite dataset we use, the steps to prepare it for using score-based diffusion, how it is adapted for the satellite forecasting task, what specific types of machine learning models are trained, and some exploration of pre-trained autoencoders. Section 4 shows the results of several score-based diffusion models trained to nowcast satellite imagery and discusses their performance compared to a couple of baseline forecasts, persistence and a mean squared error trained U-Net. Section 5 discusses lessons learned, limitations, and improvements that could be made. Lastly, Section 6 summarizes the paper.
2 Intuition for Score-Based Diffusion Models
@@ 2.1 Preface
All diffusion models discussed in this paper follow a score-based diffusion framework (EDM; Song2021; Karras2022). Before we dive into the details of the EDM framework we acknowledge another common framework for diffusion models, named Denoising Diffusion Probabilistic Models (DDPM; ho2020). We initially explored DDPMs for this project given their wide use in the literature (e.g., Asperti2024; Chen2023; Gao2023; Srivastava2024; Zhong2024), but realized the computational demand would not be viable given our hardware and the temporal demand of an operational nowcasting product. Instead, after seeing the results from Li2024, Mardani2025, Pathak2024, and Price2025, we decided to pursue score-based diffusion models which we found easier to train and evaluate. The EDM framework can be interpreted as a generalization of the DDPM approach. For more information on this topic, see ho2020 for DDPMs, see Song2021 and Karras2022 for the connection between the DDPM and EDM framework, and see references therein and several Earth and Environmental Science papers using DDPMs (Asperti2024; Chen2023; Gao2023; Srivastava2024; Zhong2024). Similarly, readers may have heard of flow matching or continuous normalizing flow (Lipman2023; Tong2024) which can be thought of as a more abstract version of diffusion models. Both define a way that maps noise to data but differ in how the network is trained and the underlying choice of formulation. For a technical description of flow matching with an application to weather forecasting see Stock2025.
@@ 2.2 Basic Idea of Diffusion Models
Generally, diffusion models can be thought of as tools to learn, and later sample from, data distributions. A diffusion model is trained on a large set of representative samples of the distribution, typically images. For example, if a diffusion model is trained on an extensive set of sample images from a specific satellite channel, the diffusion model learns to approximate this distribution. Once trained, the diffusion model can provide new samples from the distribution, i.e., it generates new (approximate) satellite imagery for that channel. This type of random but realistic image generation is called unconditional image generation. Although such ability is useful for many applications, we are more interested here in the ability to generate specific imagery, known as conditional image generation. Conditional image generation would allow, for example, to use information from a different channel of imagery or to forecast satellite imagery 10 minutes from now by conditioning on relevant imagery. In Sections 22.2 to 22.5 we focus entirely on diffusion models for unconditional imagery, while Section 22.6 discusses how to add the conditioning.
To perform image generation, diffusion models utilize a mathematically well-described diffusion process for both the training of the underlying neural network and generating new images. Namely, to learn the distribution from a given data set of images (training set), training images are partially destroyed by applying a diffusion process to the image (i.e., adding noise). This is called the “forward process”. The diffusion model is then trained to restore the structure of the images from their degenerated versions. This is called the “backwards process”, and it teaches the model to generate imagery of the same structure as the images in the training data set. In other words, the model learns to approximate the distribution represented by the training data set. Diffusion models are thus based on developing a model to reverse the diffusion equation, an approach developed mathematically over 40 years ago by anderson1982reverse. However, certain components of this approach are hard to implement in practice, and thus the full potential of this approach was made possible only once those components could be implemented through neural networks, as proposed by sohl2015deep.
@@ 2.3 Score-based Diffusion - Mathematical Foundation
Score-based diffusion is, in essence, a distribution-to-distribution transformation. We denote as $\mathcal{X}$ the distribution from which we want to generate new data (typically images, and in our case GOES IR images over time). $\mathcal{X}$ is the abstract distribution from which the training data was drawn, consisting of elements $x$. The training data is denoted as $\hat{\mathcal{X}}$ which we use as an approximation of $\mathcal{X}$. The probability density function for $\mathcal{X}$, $p(x)$, describes how representative or uncommon a particular $x$ is within the scope of $\mathcal{X}$; to make this concrete, for our example, in theory any combination of pixel values of the appropriate shape could be observed from GOES, but some combinations are far more likely than others, and that likelihood is what $p(x)$ represents.
Figure 1: Graphic illustration of what the distribution of training data looks like, both in its 2d shape as images (left) and a 1d representation for conceptualization (right, the histogram representing the pixel-value distribution).
Fig. 1 shows on the left sample images from $\hat{\mathcal{X}}$ and on the right a simplified representation of an estimation of $p(x)$ as a histogram. While $\hat{\mathcal{X}}$ is a high-dimensional distribution with complex spatial relationships, for simplicity of visualization the histogram shown is 1-dimensional and represents the distribution of the values for a single pixel (i.e., for a “1x1” image).
Figure 2: Schematic describing main task for diffusion models: seeking a way to sample from an unknown distribution, $\mathcal{X}$}, by sampling from a known distribution, such as a Gaussian distribution, $\mathcal{N}(\mu,\sigma)$.
The simplified plot of $p(x)$ in Fig. 1 notwithstanding, we have neither an explicit formulation for $p(x)$ nor knowledge of how to draw a new sample image from $\mathcal{X}$. However, we do know how to draw samples from many well-known distributions such as a Gaussian distribution $\mathcal{N}(\mu,\sigma)$ (bell-shaped curve), shown as the “known distribution” in Fig. 2. Moreover, because the Gaussian distribution is closely related to “white noise”, we can describe a process to iteratively transform images from $\hat{\mathcal{X}}$ into samples from a particular multivariate Gaussian by adding sufficient quantities of white noise; that is, we know how to map from the “unknown distribution” to the known distribution in Fig. 2. The goal of diffusion is to learn the inverse of that mapping: to learn how to map a sample from the known distribution (which is easy to generate) to a corresponding sample in the unknown distribution $\mathcal{X}$.
Figure 3: A visual diagram illustrating conceptually the forward and backward process for a diffusion model. The left column shows the simplified 1-d representations of distributions. The center column shows single image examples. The right column shows all 1-d representations of the data, which is known as $x-\sigma$ diagram in the computer science literature. Note that we switched the axes of the $x-\sigma$ diagram - in comparison to the literature - to emphasize the relationship to the histograms in the left column. Namely, each histogram on the left represents one $\sigma$ value and corresponds to one horizontal line in the $x-\sigma$ diagram on the right.
In Fig. 3, we illustrate how we can create these mappings by creating intermediate spaces with increasing amounts of noise contamination between the known and unknown distributions, which we now introduce notation for: what we have been calling $\mathcal{X}$ is now $\mathcal{X}_{0}$ (where the subscript indicates noise level 0), for each noise level $\sigma$ we have $\mathcal{X}_{\sigma}$ up to some maximum $\sigma=\sigma_{\max}$. The forward process starts with an intact image from $\hat{\mathcal{X}}$ (bottom of Fig. 3) and adds Gaussian noise repeatedly until only noise is left (top of Fig. 3, where $\mathcal{X}_{\sigma_{\max}}\approx\mathcal{N}(\mu,\sigma_{\max})$). This is straight forward and efficient (as we can combine steps to jump directly from a sample in $\mathcal{X}_{0}$ to a corresponding sample in $\mathcal{X}_{\sigma}$ for any $\sigma$ in one step), and we make use of this to expand our training set $\hat{\mathcal{X}}$ into a continuum of increasingly noisy distributions $\hat{\mathcal{X}_{\sigma}}$ (which in turn approximate the abstract distributions $\mathcal{X}_{\sigma}$). The 1D histograms for one of these spaces are shown in the left column of Fig. 3, and the complete set of 1D histograms are plotted as the horizontal slices of the rightmost plot in Fig. 3. This “$x$-$\sigma$” diagram is an illustration of the space we have created to interpolate between the known and unknown distributions.
The backwards process is the hard part and we need machine learning to do this. The goal is to turn an image of random noise (top of Fig. 3) into an image that would be within the distribution of $\mathcal{X}$} (bottom of Fig. 3), i.e., a GOES IR image in our case. The trick of the backward process is to iteratively remove the noise in the image while guiding each pixel’s value so that the modified image is representative of $\mathcal{X}$} (GOES IR image). This is an iterative process that is performed in many steps, generating a sequence of images that removes a small amount of noise at each step, which corresponds to a trajectory through the $x$-$\sigma$ space (rightmost figure in Fig. 3) that steps slowly down through the $\mathcal{X}_{\sigma}$ layers from the top (known distribution) to the bottom (unknown distribution $\mathcal{X}_{0}$). The left two columns of Fig. 3 show this process collapsed into two steps for simplicity of visualization. In each iteration, we want to move the sample $\bm{x}$ (new image that we are generating) in a direction where it will become more likely under $\mathcal{X}$ – that is, we want to maximize $p(\bm{x})$, and in particular $p(\bm{x},\sigma)$, which is the same idea of a likelihood, just specialized to a particular noise level. To know what direction to move in to maximize $p(\bm{x},\sigma)$, we use the score function, which is defined to be $\nabla_{\bm{x}}\log p(\bm{x},\sigma)$. Putting this all together, to define how to update $x$ through these iterations (i.e., as we move along this trajectory), we utilize an ordinary differential equation (ODE):
$\mathrm{d}\bm{x}=-\sigma\underbrace{\nabla_{\bm{x}}\log p\big(\bm{x};\sigma}_{\text{score function}}\big)\,\mathrm{d}\sigma.$ (1)
Figure 4: $x-\sigma$ diagram illustrating one iteration of the backward process of a diffusion model, namely the transition from one image, A, to the next image, B, that is closer to the distribution of $\mathcal{X}$, using the gradient term, $\nabla_{\bm{x}}\log p(\bm{x},\sigma)$.
Fig. 4 illustrates one step of the process, going from Image A to Image B, where we move from higher noise and lower likelihood (image A) down the gradient to less noise and higher likelihood under $\mathcal{X}$} (image B), thus closer to a GOES IR image in our case. We note that elsewhere in the diffusion literature, the reader will likely see the equivalent to Equation 1 phrased in terms of the variable $t$ and $\sigma$ as a function ($\sigma(t)$) – this is the more general form, and allows for the noise level to vary non-linearly with the variable being discretized to solve the ODE. For the full mathematical details, we refer interested readers to the foundational paper on score-based diffusion, Karras2022.